OpenStudy (anonymous):

simplify: cos(π/2-x)

OpenStudy (anonymous):

cofunction identity for this one

OpenStudy (anonymous):

\[\cos(\frac{\pi}{2}-x)=\sin(x)\] and if you draw a triangle it will be clear why this has to be true

OpenStudy (anonymous):



OpenStudy (anonymous):

got it thank you

OpenStudy (anonymous):

is there any other way i can do this without the diagram?